OpenStudy (anonymous):

X^2 – 2x – 35 = 0 Solve for x, show work.

OpenStudy (anonymous):

(x+7)(x-5)=0 x=-7,5

OpenStudy (anonymous):

7 times 5 = 35 -7 + 5 gives you -2 so x^2-2x-35 = (x-7)(x+5)=0 then solve for x-7=0, and x+5 = 0, -5 and 7 are your zeroes.